Solve the given initial value problem using the method of Laplace transforms. 26, Osts 7 y + 2y + 20y = g(t), y(0)= -5, y(

Identify p(t), q(t) and r(t).
Answer #1

In unik step function notation, get = 26 (ult-ol-ult- +-71) +52cuit-z-ULE-14) 26 ult) +264(7-77-52417-14) Ligety=:26 Lsulty +7 sta Now, Il strs+26 (sti )t =? (A+127 25 =et 12 sh $7.25 [:22 F (stail = eat & F (S) et [El htt) 15 =ét (co cos (st) + Siti sut 21426) 22 to } St 2 st2stro 26-to [e *(cos(56+ + usi sing sin (st) 5 using and and shifting throrem, -74 26 e list 25+5 1 ) (t-7) skin (544-7) 466-73 yet set(cos (58) + 1 Sin (SH) Hi-tet(cop (57tús tin (su)] +(1-lett (cos (5 (+-71) + ult-ly)